Find the work done by the force field F on a particle that moves along the curve C.

    F(x,y)=xy  i+x^2  j
 C: x=y2 from (0,0) to (4,2)



Enter the exact answer as an improper fraction, if necessary.

W=
Answer #1

the Curve On а + • It is asked to find the work done by the force field Paxy + x² j on the particle theet moves along C: 2=y?dets parameterize the wine C: x=y² from (0,0) to (4, 2) as fellenes og zit gi 2 y=t x=y²=t 7 (t) = t² i + tj y = t yo » whenat t² + + x1 = 8th rocha 3th at ..We 2 Note ) Saranx= ahhh te t=o 5 2 na- 3.t 5 (19-05)=*32 - 96 W = 96 5